OpenStudy (anonymous):

-8(4x) simplify and justify

OpenStudy (anonymous):

you would just multiply them

OpenStudy (anonymous):

I don't know what they want with justify though

OpenStudy (anonymous):

-32x. the justification is that you distribute the variables through the parentheses like a boss

OpenStudy (anonymous):

@moser90 The justification is the associative property of multiplication. \[a(b\cdot c) = a\cdot b \cdot c\]There is an associative property of addition too:\[a+(b + c) = a +b + c\]

OpenStudy (anonymous):

thank you I learn something too I have a hard time with that